OpenStudy (anonymous):

Lauren will make annual contributions in the amount of $4,770, on average, to a 401(k) over the next 34 years. Her employer will match 60% of her contributions. She is currently being taxed at 33%, but anticipates being taxed at 15% upon retirement. If her account grows at an average rate of 5.5% annually, what is the value of Lauren’s 401(k) upon retirement? $381,435.69 $395,228.47 $481,057.72 $610,297.11

OpenStudy (anonymous):

5.5% annually

OpenStudy (anonymous):

$4,770

OpenStudy (anonymous):

34 years.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

FVOA=4770*((1+0.055)^(1*34)-1))/(0.055) ?

OpenStudy (anonymous):

then 5.5% - 4770 - 34 ?

OpenStudy (anonymous):

@turtleluver

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

is this the equation for 'uniform series payments'? FVOA=4770*((1+0.055)^(1*34)-1))/(0.055) ?

OpenStudy (anonymous):

if so, then you are doing it correctly ^_^

OpenStudy (anonymous):

@DemolisionWolf I got 448747.87 what do I do next o.o

OpenStudy (anonymous):

@Dogpile so ur equatoin is the one for uniform series then? not single payment or gradient series

OpenStudy (anonymous):

I think soo I dont really know o.O

OpenStudy (anonymous):

it's been a while since i have used that equatoin lemme look..

OpenStudy (anonymous):

ok :3

OpenStudy (anonymous):

this is the same one right? \[F=A \frac{ (1+i)^n - 1 }{ i}\] [F=A \frac{ (1+i)^n - 1 }{ i}\]

OpenStudy (anonymous):

o.o no thats a different formula

OpenStudy (anonymous):

like this right? (its the same) \[FVOA=4770\frac {((1+0.055)^{1*34}-1))}{(0.055)}\]

OpenStudy (anonymous):

yaya, the second one

OpenStudy (anonymous):

so, we have to fix the A term you have. the phrase, " Her employer will match 60% of her contributions." means A = 4770 + 4770(.6)

OpenStudy (anonymous):

how does this play into our equation? i'm a bit hesitant about it, "She is currently being taxed at 33%, but anticipates being taxed at 15% upon retirement. "

OpenStudy (anonymous):

I think your suppose to ignore because questions like these tend to want to throw you off :/

OpenStudy (anonymous):

i'm with you on this one! because also, the money she is putting in every year should have already been taxed, and the 4770 is what is remaining. so now we have: \[F=(4770+(.6)4770) \frac{ (1+0.055)^{34} - 1 }{ 0.055}\] **[F=(4770+(.6)4770) \frac{ (1+0.055)^{34} - 1 }{ 0.055}\]

OpenStudy (anonymous):

whad u get for FVOA, or as I call it, F?

OpenStudy (anonymous):

thats what I get too. ^_^

OpenStudy (anonymous):

What we do next ?

OpenStudy (anonymous):

FVOA = 71,7996.59 and we have: $381,435.69 $395,228.47 $481,057.72 $610,297.11 to pick from... what did we do wrong... I wonder if the 'tax' part needs to be in this

OpenStudy (anonymous):

I think sooo you know how to put them in the problem ?

OpenStudy (anonymous):

gosh, I can't figure this one out..